# Gandalf the Red: Adaptive Security for LLMs

### v1

@@ Abstract
Current evaluations of defenses against prompt attacks in large language model (LLM) applications often overlook two critical factors: the dynamic nature of adversarial behavior and the usability penalties imposed on legitimate users by restrictive defenses. We propose D-SEC (Dynamic Security Utility Threat Model), which explicitly separates attackers from legitimate users, models multi-step interactions, and rigorously expresses the security-utility in an optimizable form. We further address the shortcomings in existing evaluations by introducing Gandalf, a crowd-sourced, gamified red-teaming platform designed to generate realistic, adaptive attack datasets. Using Gandalf, we collect and release a dataset of 279k prompt attacks. Complemented by benign user data, our analysis reveals the interplay between security and utility, showing that defenses integrated in the LLM (e.g., system prompts) can degrade usability even without blocking requests. We demonstrate that restricted application domains, defense-in-depth, and adaptive defenses are effective strategies for building secure and useful LLM applications. Code is available at https://github.com/lakeraai/dsec-gandalf.
@@ 1  Introduction
Defending against prompt attacks is challenging due to the ability of LLMs to interpret natural language, multimodal inputs, code, and obfuscated commands, making it easy to hide malicious commands within benign-looking inputs. In the literature, many approaches for defending against prompt attacks have been proposed. They can be roughly divided into three categories: (i) Input/output classifier defenses that use the input and/or the resulting LLM output to classify whether a given prompt is malicious (e.g., Kumar et al., 2024; Ayub and Majumdar, 2024; Kim et al., 2024; Sawtell et al., 2024), (ii) LLM internal defenses that modify the LLM itself either through the system prompt, fine-tuning or other post-training methods, to ensure it adheres only to the original task (e.g., Bai et al., 2022; Wei et al., 2023; Chen et al., 2024a, b; Wallace et al., 2024; Zhou et al., 2024; Piet et al., 2024), and (iii) prompt modification or sanitization defenses that restructure the prompt or remove any potentially malicious parts of the input (e.g., Learn Prompting, 2024; Hines et al., 2024; Chen et al., 2024a). None of the existing defenses is perfect (Liu et al., 2024b), and all of them can impact the utility of the LLM application by falsely blocking legitimate user requests. Since defenses can be embedded within the LLM itself (e.g., via the system prompt), they can significantly impact user utility by impacting the behavior of the application. Evaluating the security of an LLM application should, therefore, also account for user utility, a factor often overlooked by most proposed approaches. There are essentially two approaches for evaluating the security of a defense.
The first approach is to use publicly available prompt attack benchmarks (e.g., Yi et al., 2023; Liu et al., 2024c; Debenedetti et al., 2024; Mazeika et al., 2024; Chao et al., 2024; Zhu et al., 2024) to assess how many attacks are blocked by the defenses. Many attacks in these benchmarks are misleading as they are non-adaptive and easily blocked by model providers, resulting in an overly optimistic evaluation.
@@
The second approach is to use red-teaming to probe the LLM application. This can either be done manually or automatically using, for instance, a second LLM to optimize attacks (e.g., Perez et al., 2022; Ganguli et al., 2022; Deng et al., 2023). The resulting attacks are generally more realistic as they are constructed adaptively. However, manual red-teaming is expensive, and current automatic approaches often lead to less diverse attacks. Both benchmark and red-teaming approaches suffer from the difficulty of assessing whether an attack was truly successful. To circumvent this, evaluations often focus on detecting malicious intent instead. However, this focus on intent is limited and can lead to overly optimistic results, as defenses might stop obvious failures while missing subtle yet successful attacks.
@@
In this work, we propose the Dynamic Security Utility Threat Model (D-SEC), which (i) captures the adaptive nature of attacks and defenses and (ii) balances security and utility by separately modeling attackers and benign users. This model operates in sessions consisting of multiple individual transactions (either attacks or legitimate user prompts), each of which can be influenced by feedback from prior transactions. This provides practitioners with a principled way to evaluate defenses for LLM applications and explicitly optimize the defense to achieve the desired security-utility trade-off.
@@
To address the limitations of existing evaluation approaches, we propose a crowd-sourced red-teaming platform (Schulhoff et al., 2023; Toyer et al., 2023), called Gandalf, [^2] https://gandalf.lakera.ai which gamifies the process of generating prompt attacks and can be used to generate realistic and diverse adaptive attacks. In Gandalf, attacks are automatically labeled based on the success indicator given by whether the password was extracted rather than intent. Using Gandalf, we constructed a large-scale, high-quality attack dataset, which we make publicly available.333 https://huggingface.co/datasets/Lakera/gandalf-rct We then perform an extensive analysis to gain generalizable insights into promising defense strategies for D-SEC.
@@
Our contributions are fourfold: (i) We introduce D-SEC and corresponding metrics that provide a principled and comprehensive way of evaluating defenses for LLM applications. (ii) We present the crowd-sourced red-teaming platform Gandalf and release a large-scale dataset of diverse attacks. (iii) We provide empirical evidence for the importance of the security-utility trade-off in LLM applications. (iv) We show empirically that three defense strategies allow us to design defenses that optimize the security-utility trade-off: restricting the application domain, aggregating multiple defenses in a defense-in-depth approach, and using adaptive defenses.
@@ 2  Dynamic Security-Utility
2.1  A Primer on Prompt Attacks
We focus on scenarios where attackers target LLM applications by crafting prompts that manipulate the LLM to deviate from its intended behavior. We call any such attack a prompt attack. Prompt attacks exploit the fact that LLMs do not have a strict, inherently enforced functional separation between developer commands used to define the intended behavior (e.g., system prompts) and user or other external inputs (e.g., documents or websites in RAG applications). Similar to SQL injection attacks (Halfond et al., 2006), this allows attackers to hide malicious commands, ranging from direct instructions to subtle manipulation patterns within any text that is passed as input to the LLM. The literature typically distinguishes between direct prompt attacks (e.g., jailbreaks), where the attacker sends inputs directly to the LLM, and indirect prompt attacks, where the attacker modifies external documents that the LLM accesses.
@@ 2.2  Threat Model
Figure 1: Dynamic Security Utility Threat Model (D-SEC). In this threat model, there are three actors: the developer, the attacker, and the user. The developer creates an LLM application $\mathcal{M}$ and defends it with a defense $\mathcal{D}$. Users and attackers interact with $\mathcal{M}$ in sessions, which are sequences of transactions that consist of the user/attacker sending a prompt/attack and receiving feedback that can be used in the next transaction. A full sequence of transactions is called a session. The goal of the developer is to design $\mathcal{D}$ in a way that defends against the attacker but only minimally impacts the utility of the user.
@@
D-SEC involves three actors: the developer, the user, and the attacker. The system under consideration consists both of an unprotected LLM application $\mathcal{M}$ and a defense (or security layer) $\mathcal{D}$. The developer’s goal is to design $\mathcal{D}$ such that it prevents attackers from exploiting $\mathcal{M}$, while not negatively impacting the user. We consider $\mathcal{M}$ to be a simple LLM application in this paper, but it could also be an LLM agent that has access to a library of tools. The defense $\mathcal{D}$ can consist of a combination of input defenses that check the input for maliciousness, internal defenses that are directly embedded in the LLM (e.g., via the system prompt, and fine-tuning), and output defenses that use the input together with the output to check whether there was an attack. In conversational LLM applications, input and output defenses generally block the output by replacing it with a refusal message if an attack is detected. In contrast, internal defenses modify the LLM itself to produce secure outputs, which can often make it difficult to detect whether a defense has even detected an attack at all. Internal defenses are unique to LLM applications and do not exist in traditional software security layers. While they can provide strong security, they can also lead to undesirable impacts on the usability of the application, see Section 4.1. In agentic systems, however, this distinction becomes less relevant, as both internal and external defenses can significantly disrupt execution flow, resulting in pronounced impacts on usability.
D-SEC models attackers and users separately. Both interact with $\mathcal{M}$ in sessions, which are sequences of transactions (of varying length) that consist of a submitted prompt (called an attack in the case of an attacker) and the resulting feedback from $\mathcal{M}$ under defense $\mathcal{D}$. We assume the attacker sessions are drawn from an attacker distribution $P_{\mathcal{M},\mathcal{D}}^{A}$ and the user sessions from a user distribution $P_{\mathcal{M},\mathcal{D}}^{U}$. Section 4.1 shows the importance of selecting a user distribution relevant to the application. Since there are multiple transactions per session, the attacker can use the feedback as a reward to optimize the attacks in subsequent transactions. Viewing the threat model as dynamic not only provides a more accurate measure of security but also suggests defensive strategies that take into account the dynamic nature of the problem (see Section 4.2). Here, we focus on the black-box setting where the feedback consists of the LLM response or a refusal message from $\mathcal{D}$, but D-SEC also applies to gray- or white-box settings by including additional information in the feedback, such as gradients.
@@ 2.4  Metrics to Evaluate Security and Utility
The attacker failure rate is equal, in the case of single transaction sessions, to one minus the attack success rate that is sometimes used in the literature (e.g., Chao et al., 2024). AFR can be estimated by
@@
SCR focuses exclusively on whether transactions are blocked. If the defenses are internal (i.e., part of the LLM itself), this can be misleading since it is hard to detect blocks (in our analysis, we use a simple classifier to detect refusals, see Appendix C.2) and the application responses might degrade without being blocked. In such cases, other metrics might be preferable, e.g., the relative decrease in length of an LLM response between the undefended and defended LLM (see Section 4.1). More generally, D-SEC can also be extended to include other utility dimensions, such as the cost of running the application.
In this work, we consider all exploits to have the same severity. In practice, one might adapt the metrics to weigh attacks by their severity and likelihood of being exploited, emphasizing high-impact vulnerabilities that are more likely to occur.
@@ 2.5  Maximizing Developer Utility
The critical question for the developer is how to design a defense that maximizes the developer utility $V^{\lambda}(\mathcal{D})$. We found that three non-exclusive defense strategies are particularly promising when constructing defenses for LLM applications (see Section 4.2): (i) Restricting the application domain so that the LLM has a more narrow use-case which makes it easier to defend. (ii) Employing a defense-in-depth strategy (Mughal, 2018) that combines multiple, ideally unrelated types of defenses together. (iii) Using adaptive defenses that take into account previous transactions in the same session to better detect attackers. For each procedure, the developer can optimize the defense in different ways to achieve maximal developer utility:
For (i), we can select how much and in what way to restrict the application domain. For example, we could start with various system prompts that define the application domain in increasingly restrictive ways, asking the LLM to refuse requests that are not relevant to the application domain. We can then evaluate the developer utility for each and select the one that maximizes it. As we show in Section 4.2, a more restrictive application domain generally increases security.
For (ii), we can adjust how the individual defenses are aggregated. For example, aggregating multiple defenses using an ‘or’-aggregation (i.e., blocking a transaction if at least one of the defenses blocks it) increases security but likely also reduces utility. How exactly security and utility are affected by different aggregations depends on the interaction between the defenses. In practice, we propose to optimize over the aggregations to achieve the optimal security-utility trade-off. To make this more concrete, let $\mathcal{D}_{1},\ldots,\mathcal{D}_{K}$ denote a sequence of non-adaptive defenses, $f:\{0,1\}^{K}\rightarrow\{0,1\}$ an arbitrary aggregation function and $\mathcal{D}_{f}$ the defense that blocks a transaction $W$ if and only if $f(\mathds{1}(\mathcal{D}_{1}\text{ blocks }W),\ldots,\mathds{1}(\mathcal{D}_{K}\text{ blocks }W))=1$. We then propose to use the aggregation function
@@
where we assume for simplicity that the maximizer is unique. Hence, the optimal aggregation can be selected by an exhaustive search (if $K$ is sufficiently small). We illustrate how this is done in practice in Section 4.2.
For (iii), there are several ways in which defenses can be made adaptive. For simplicity, we focus on the most basic adaptive defenses that limit the number of times a defense flags a potential attack (i.e., blocks a transaction) to $T$ before blocking the entire session such that the attacker can no longer submit prompts. For such defenses, there is a clear security-utility trade-off: Smaller $T$ increase security but decrease utility, and larger $T$ increase utility but decrease security. As we show in Section 4.2, one can select an optimal threshold that maximizes the developer utility.
@@ 3  Gandalf
A crucial component of D-SEC is that attacks can be adaptive, which cannot be captured by existing static evaluation methods. To account for this shortcoming, manual and automatic red-teaming, in which application-defense combinations are probed with sequentially optimized attacks, have been proposed. We believe that high-quality red-teaming provides the gold standard for evaluating the security of a defense. Nevertheless, ensuring a diverse and accurate set of attacks is challenging: manual red-teaming is laborious, and automated LLM-based approaches struggle with diversity and label accuracy (e.g. because attacks contain obfuscations to avoid detection).
We propose a red-teaming approach based on crowd-sourcing called Gandalf. Gandalf is a web application that turns prompt attacks against specific defenses into a publicly accessible interactive game, which addresses these challenges. The idea of the game is simple: an application LLM is set up with a use-case (e.g., “Be a friendly AI assistant.”) and told a password in its system prompt that the user needs to extract by interacting with the application LLM. Each level corresponds to different types and strengths of defenses that are added to secure the password. If the player is able to guess the password, they move on to the next level (see Figure 2). The game incentivizes the player to construct creative attacks while simultaneously labeling whether the attack was successful by requiring the password to move on to the next level. From the perspective of D-SEC, each player playing a single level corresponds to an attacker session and can be seen as a random draw from $P_{\mathcal{D}}^{A}$, where $\mathcal{D}$ is the level defense.
Gandalf has been online since May 2023 and so far more than one million players have submitted over 40 million prompts and password guesses. Gandalf users have spent more than a combined 25 years playing the game. The value of this data lies in the diversity of the attacks and the realistic adaptive nature under which they are constructed. As we show below, it can be used to gain valuable insights into which defense strategies provide good security in D-SEC.

### v2

@@ Abstract
Current evaluations of defenses against prompt attacks in large language model (LLM) applications often overlook two critical factors: the dynamic nature of adversarial behavior and the usability penalties imposed on legitimate users by restrictive defenses. We propose D-SEC (Dynamic Security Utility Threat Model), which explicitly separates attackers from legitimate users, models multi-step interactions, and expresses the security-utility in an optimizable form. We further address the shortcomings in existing evaluations by introducing Gandalf, a crowd-sourced, gamified red-teaming platform designed to generate realistic, adaptive attack. Using Gandalf, we collect and release a dataset of 279k prompt attacks. Complemented by benign user data, our analysis reveals the interplay between security and utility, showing that defenses integrated in the LLM (e.g., system prompts) can degrade usability even without blocking requests. We demonstrate that restricted application domains, defense-in-depth, and adaptive defenses are effective strategies for building secure and useful LLM applications.
@@ 1  Introduction
Defending against prompt attacks is challenging due to the ability of LLMs to interpret natural language, multimodal inputs, code, and obfuscated commands, making it easy to hide malicious commands within benign-looking inputs. In the literature, many approaches for defending against prompt attacks have been proposed. They can be roughly divided into three categories: (i) Input/output classifier defenses that use the input and/or the resulting LLM output to classify whether a given prompt is malicious (e.g., Kumar et al., 2024; Ayub and Majumdar, 2024; Kim et al., 2024; Sawtell et al., 2024), (ii) LLM internal defenses that modify the LLM itself either through the system prompt, fine-tuning or other post-training methods, to ensure it adheres only to the original task (e.g., Bai et al., 2022; Wei et al., 2023; Chen et al., 2024a, b; Wallace et al., 2024; Zhou et al., 2024; Piet et al., 2024), and (iii) prompt modification or sanitization defenses that restructure the prompt or remove any potentially malicious parts of the input (e.g., Learn Prompting, 2024; Hines et al., 2024; Chen et al., 2024a). None of the existing defenses is perfect (Liu et al., 2024b), and all of them can impact the utility of the LLM application by falsely blocking legitimate user requests. Since defenses can be embedded within the LLM itself (e.g., via the system prompt), they can significantly impact user utility by impacting the behavior of the application. Evaluating the security of an LLM application should, therefore, also account for user utility, a factor overlooked by most proposed approaches. There are essentially two approaches for evaluating the security of a defense.
@@
The first approach is to use publicly available prompt attack benchmarks (e.g., Yi et al., 2023; Liu et al., 2024c; Debenedetti et al., 2024; Mazeika et al., 2024; Chao et al., 2024; Zhu et al., 2024) to assess how many attacks are blocked by the defenses. Many attacks in these benchmarks are misleading as they are non-adaptive and easily blocked by model providers, resulting in an overly optimistic evaluation. The second approach is to use red-teaming to probe the LLM application. This can either be done manually or automatically, using, for instance, a second LLM to optimize attacks (e.g., Perez et al., 2022; Ganguli et al., 2022; Deng et al., 2023). The resulting attacks are generally more realistic as they are constructed adaptively. However, manual red-teaming is expensive, and current automatic approaches often lead to less diverse attacks. Both benchmark and red-teaming approaches suffer from the difficulty of assessing whether an attack was truly successful. To circumvent this, evaluations often focus on detecting malicious intent instead. However, this focus on intent is limited and can lead to overly optimistic results, as defenses might stop obvious failures while missing subtle yet successful attacks.
@@
In this work, we propose the Dynamic Security Utility Threat Model (D-SEC), which (i) captures the adaptive nature of attacks and defenses and (ii) balances security and utility by separately modeling attackers and benign users. This model operates in sessions consisting of multiple transactions (either attacks or legitimate user prompts), each of which can be influenced by feedback from prior transactions. This provides a principled way to evaluate defenses for LLM applications and explicitly optimize the defense to achieve the desired security-utility trade-off.
@@
To address the limitations of existing evaluation approaches, we propose a crowd-sourced red-teaming platform (Schulhoff et al., 2023; Toyer et al., 2023), called Gandalf, [^2] https://gandalf.lakera.ai which gamifies the process of generating realistic and diverse adaptive prompt attacks. In Gandalf, attacks are automatically labeled based on the success indicator given by whether the password was extracted rather than intent. Using Gandalf, we constructed a large-scale, high-quality attack dataset, which we make publicly available. [^3] https://huggingface.co/datasets/Lakera/gandalf-rct We then perform an extensive analysis to gain generalizable insights into promising defense strategies for D-SEC.
@@
Our contributions are fourfold: (i) We introduce D-SEC and corresponding metrics that provide a principled and comprehensive way of evaluating defenses for LLM applications. (ii) We present the crowd-sourced red-teaming platform Gandalf and release a large-scale dataset of diverse attacks. (iii) We provide empirical evidence for the importance of the security-utility trade-off. (iv) We show empirically that three defense strategies allow us to design defenses that optimize the security-utility trade-off: restricting the application domain, aggregating multiple defenses in a defense-in-depth approach, and using adaptive defenses.
@@ 2  Dynamic Security-Utility
We focus on scenarios where attackers target LLM applications by crafting prompts that manipulate the LLM to deviate from its intended behavior. We call any such attack a prompt attack. Prompt attacks exploit the fact that LLMs do not have a strict, inherently enforced functional separation between developer commands used to define the intended behavior (e.g., system prompts) and user or other external inputs (e.g., documents or websites in RAG applications). Similar to SQL injection attacks (Halfond et al., 2006), this allows attackers to hide malicious commands, ranging from direct instructions to subtle manipulation patterns within any text that is passed as input to the LLM. The literature typically distinguishes between direct prompt attacks (e.g., jailbreaks), where the attacker sends inputs directly to the LLM, and indirect prompt attacks, where the attacker modifies external sources that the LLM accesses.
@@ 2.1  Threat Model
Figure 1: Dynamic Security Utility Threat Model (D-SEC). In this threat model, there are three actors: the developer, the attacker, and the user. The developer creates an LLM application $\mathcal{M}$ and defends it with a defense $\mathcal{D}$. Users and attackers interact with $\mathcal{M}$ in sessions, which are sequences of transactions that consist of the user/attacker sending a prompt/attack and receiving feedback that can be used in the next transaction. A full sequence of transactions is called a session. The goal of the developer is to design $\mathcal{D}$ to defend against the attacker but only minimally impact user utility.
@@
D-SEC involves three actors: the developer, the user, and the attacker. The system under consideration consists both of an unprotected LLM application $\mathcal{M}$ and a defense (or security layer) $\mathcal{D}$. The developer’s goal is to design $\mathcal{D}$ such that it prevents attackers from exploiting $\mathcal{M}$, while not negatively impacting the user. We consider $\mathcal{M}$ to be a simple LLM application in this paper, but it could also be an LLM agent that has access to a library of tools. The defense $\mathcal{D}$ can consist of a combination of input defenses that check the input for maliciousness, internal defenses that are directly embedded in the LLM (e.g., via the system prompt, and fine-tuning), and output defenses that use the input together with the output to check whether there was an attack. In conversational LLM applications, input and output defenses generally block the output by replacing it with a refusal message if an attack is detected. In contrast, internal defenses modify the LLM itself to produce secure outputs, which can make it difficult to determine whether a defense has detected an attack. Internal defenses are unique to LLM applications and do not exist in traditional software security. While they can provide strong security, they can impact the usability of the application, see Section 4.1. In agentic systems, however, this distinction becomes less relevant, as both internal and external defenses can significantly disrupt execution flow, resulting in pronounced impacts on usability.
D-SEC models attackers and users separately. Both interact with $\mathcal{M}$ in sessions, which are sequences of transactions (of varying length) that consist of a submitted prompt (called an attack in the case of an attacker) and the resulting feedback from $\mathcal{M}$ under defense $\mathcal{D}$. We assume the attacker sessions are drawn from an attacker distribution $P_{\mathcal{M},\mathcal{D}}^{A}$ and the user sessions from a user distribution $P_{\mathcal{M},\mathcal{D}}^{U}$. Section 4.1 shows the importance of selecting a user distribution relevant to the application. Since there are multiple transactions per session, the attacker can use the feedback as a reward to optimize subsequent attacks. Viewing the threat model as dynamic not only provides a more accurate measure of security but also suggests defensive strategies that take into account the dynamic nature of the problem (see Section 4.2). While we focus on the black-box setting where feedback consists of the LLM response or a refusal message from $\mathcal{D}$, D-SEC also applies to gray- or white-box settings by adding further information to the feedback, e.g., gradients.
@@ 2.3  Metrics to Evaluate Security and Utility
The attacker failure rate is equal, for single transaction sessions, to one minus the attack success rate that is used in the literature (e.g., Chao et al., 2024). AFR can be estimated by
@@
SCR focuses exclusively on whether transactions are blocked. If the defenses are internal (i.e., part of the LLM itself), they can be misleading, as blocks are hard to detect (we use a classifier to identify refusals; see Appendix C.2), and responses may degrade without being blocked. In such cases, alternative metrics, like the relative decrease in response length between undefended and defended LLMs, may be more suitable (see Section 4.1). More generally, D-SEC can also be extended to include other utility dimensions, such as the cost of running the application.
In this work, we consider all exploits to have the same severity. In practice, one might adapt the metrics to weigh attacks by their severity and likelihood of being exploited, emphasizing high-impact vulnerabilities more likely to occur.
@@ 2.4  Maximizing Developer Utility
The critical question for the developer is how to design a defense that maximizes the developer utility $V^{\lambda}(\mathcal{D})$. We found that three non-exclusive defense strategies are promising (see Section 4.2): (i) Restricting the application domain so that the LLM has a more narrow use-case which makes it easier to defend. (ii) Employing a defense-in-depth strategy (Mughal, 2018) that combines multiple, ideally unrelated types of defenses together. (iii) Using adaptive defenses that take into account previous transactions in the same session to better detect attackers.
For (i), we can vary the degree and type of restriction, for example by using increasingly restrictive system prompts that make the LLM reject off-topic requests. We then measure the developer utility of each option and select the one that maximizes it. As shown in Section 4.2, more restrictive domains generally enhance security.
For (ii), we can adjust how the individual defenses are aggregated. For example, aggregating multiple defenses using an ‘or’-aggregation (i.e., blocking a transaction if at least one of the defenses blocks it) increases security but likely reduces utility. How exactly security and utility are affected by different aggregations depends on the interaction between the defenses. In practice, we propose to optimize over the aggregations to achieve the optimal security-utility trade-off. Let $\mathcal{D}_{1},\ldots,\mathcal{D}_{K}$ denote a sequence of non-adaptive defenses, $f:\{0,1\}^{K}\rightarrow\{0,1\}$ an arbitrary aggregation function and $\mathcal{D}_{f}$ the defense that blocks a transaction $W$ if and only if $f(\mathds{1}(\mathcal{D}_{1}\text{ blocks }W),\ldots,\mathds{1}(\mathcal{D}_{K}\text{ blocks }W))=1$. We then propose to use the aggregation function
@@
where we assume for simplicity that the maximizer is unique. The optimal aggregation, if $K$ is sufficiently small, can be selected by an exhaustive search, as shown in Section 4.2.
For (iii), there are several ways in which defenses can be made adaptive. For simplicity, we focus on simple defenses that limit the number of times a defense flags a potential attack (i.e., blocks a transaction) to $T$ before blocking the entire session such that the attacker can no longer submit prompts. There is a clear security-utility trade-off: Smaller $T$ increase security but decrease utility, and vice-versa for larger $T$. We show in Section 4.2 how to select an optimal threshold to maximize developer utility.
@@ 3  Gandalf
A crucial component of D-SEC is that attacks can be adaptive, which cannot be captured by existing static evaluation methods. To account for this shortcoming, manual and automatic red-teaming, in which application-defense combinations are probed with sequentially optimized attacks, have been proposed. We believe high-quality red-teaming is the gold standard for evaluating defense security. Nevertheless, ensuring a diverse and accurate set of attacks is challenging: manual red-teaming is laborious, and automated LLM-based approaches struggle with diversity and label accuracy (e.g. because attacks contain obfuscations to avoid detection).
To address this, we propose Gandalf, a red-teaming approach based on crowd-sourcing. Gandalf is a web platform that turns prompt attacks against LLM defenses into a publicly accessible interactive game. The game’s concept is simple: an application LLM is set up with a use-case (e.g., “Be a friendly AI assistant.”) and told a password in its system prompt that the user needs to extract. Each level corresponds to different types and strengths of defenses that are added to secure the password. If the player is able to guess the password, they move on to the next level (see Figure 2). The game incentivizes players to create creative attacks while simultaneously labeling whether the attack was successful by requiring the password to move on to the next level. In terms of D-SEC, each player playing a level corresponds to an attacker session and can be seen as a random draw from $P_{\mathcal{D}}^{A}$, where $\mathcal{D}$ is the level defense.
Gandalf has been online since May 2023 and so far more than one million players have submitted over 40 million prompts and password guesses. Gandalf users have spent more than a combined 25 years playing the game. The value of this data lies in the diversity of the attacks and the realistic adaptive nature under which they are constructed. As we show below, it provides valuable insights into which defense strategies lead to good security in D-SEC.